OpenStudy (anonymous):

The radian measure equivalent to -120º is ______.

jimthompson5910 (jim_thompson5910):

Multiply by pi/180 to convert a degree angle to a radian angle

OpenStudy (mathstudent55):

All around a circle is 360 deg and 2 pi rad, so we have \(360^o = 2 \pi ~rad\) Divide both sides by 2 to get \(180^o = \pi ~rad\) Now divide both sides by 180 deg to get: \( 1 = \dfrac{\pi ~rad}{180^o} \) That last fraction is equal to 1. You know that multiplying a number by 1 does not change the number (this is a property of multiplication). That means you can multiply \(120^o\) by the fraction without changing the number. You're only changing the form. That is how you convert degrees to radians. \(120^o \times \dfrac{\pi ~rad}{180^o} =\)